OpenStudy (just_one_last_goodbye):

hey Guys I have a DBA and I need help on just one small thing I dont understand before I call my teacher, 2x - y -3w x=4 y=5 w=-6. I would Also like to know whats the name of the algebra operation. Thank you.

OpenStudy (teddyiswatshecallsme):

What are you solving for? what are the instructions?

Directrix (directrix):

I don't understand what the task is. Is it to evaluate the expression ( 2x - y -3w ) for x=4 y=5 w=-6 by substitution.

Directrix (directrix):

Substitution may be the word of what you are thinking.

OpenStudy (just_one_last_goodbye):

Yes @Directrix thats its evaluate

OpenStudy (teddyiswatshecallsme):

Okay...

OpenStudy (just_one_last_goodbye):

I failed my first attempt on the DBA cause i got this one wrong I would like to know how to solve it ^_^

OpenStudy (texaschic101):

2x - y - 3w 2(4) - (5) - 3(-6) so whats 2 * 4 =

OpenStudy (just_one_last_goodbye):

8

OpenStudy (anonymous):

FVLS?

OpenStudy (texaschic101):

8 ...then we have -5.....now whats 3 * -6

OpenStudy (just_one_last_goodbye):

Connexus academy but same thing just different names

OpenStudy (just_one_last_goodbye):

@texaschic101 so if the number is by its self its automatically a negative 5?

OpenStudy (anonymous):

\[Cool :)\]

OpenStudy (just_one_last_goodbye):

im guessing 3* -6= -18

OpenStudy (texaschic101):

2(4) - (5) - 3(-6) = 8 - 5 - (-18) --- a negative times a negative is a positive 8 - 5 + 18 ....now just finish with the simple arithmetic and what do you get ?

OpenStudy (texaschic101):

can you finish ?

OpenStudy (just_one_last_goodbye):

so since subtraction comes first its... 8-5 = 3 and 3 + 18=21

OpenStudy (texaschic101):

correct....answer is 21

OpenStudy (just_one_last_goodbye):

hmm I feel stupid now xD seems so easy

OpenStudy (texaschic101):

don't feel stupid...you probably had trouble with the signs....just do it enough and all this will come natural to you

OpenStudy (just_one_last_goodbye):

thanks @texaschic101 oh btw idk if u remember me but im UsArmy3947

OpenStudy (texaschic101):

thats you....cool....good to see you again :)

OpenStudy (just_one_last_goodbye):

nice to see you 2 ^_^

OpenStudy (just_one_last_goodbye):

omg im so happy now thanks so much @texaschic101 cant wait till my teacher calls back

OpenStudy (texaschic101):

tell you you know exactly what your doing and can ace any test given...lol

OpenStudy (just_one_last_goodbye):

@texaschic101 do you know whats the name of this operation or no?